OpenStudy (thomas5267):

Prove that 2 linearly independent vectors in \(\mathbb{R}^2\) is a basis for \(\mathbb{R}^2\).

ganeshie8 (ganeshie8):

I think you need to show the vectors span \(\mathbb R^2\)

OpenStudy (thomas5267):

How about this: \[ \mathbf{u}= \begin{bmatrix} u_1\\ u_2 \end{bmatrix} , \, \mathbf{v}= \begin{bmatrix} v_1\\ v_2 \end{bmatrix} ,\, \mathbf{b}= \begin{bmatrix} b_1\\ b_2 \end{bmatrix}\\ \text{We have to prove that for all }\mathbf{b}\in\mathbb{R}^2, \begin{bmatrix} \mathbf{u}&\mathbf{v} \end{bmatrix} \begin{bmatrix} x_1\\ x_2 \end{bmatrix} =\mathbf{b} \text{ has a unique solution.}\\ \] The matrix \(\begin{bmatrix}\mathbf{u}&\mathbf{v}\end{bmatrix}\) has full rank. Therefore, it is invertible. Thus, the equation has a unique solution. Am I doing circular reasoning?

ganeshie8 (ganeshie8):

i think so, the statement we're trying to prove is equivalent to below : prove that a basis of "n" dimensional vector space contains exactly "n" vectors

ganeshie8 (ganeshie8):

which is equivalent to proving : prove that the vectors in a full rank nxn matrix span the n dimensional space

ganeshie8 (ganeshie8):

so i don't think we can use rank here..

OpenStudy (thomas5267):

But the definition of dimension of a vector space is the number of vectors in the basis of that vector space...

ganeshie8 (ganeshie8):

yes, so is it satisfying to conclude like this : since basis of n dimensional vector space contains 2 independent vectors, the 2 given independent vectors represent a basis. QED. ? i thought we need to say something about why "n" given vectors represent a basis for "n" dimensional vector space